OpenStudy (anonymous):

find the point on the curve y^2 = 4x which is nearest to the point (2,-8) find the point on the curve y^2 = 4x which is nearest to the point (2,-8) @Mathematics

OpenStudy (zarkon):

us the distance formula (then take the derivative...set =0...yada yada yada)

OpenStudy (anonymous):

can't understand can you please explain with the steps!!!!

OpenStudy (zarkon):

plug your function and equation into the distance formula \[D^2=(x_2-x_1)^2+(y_2-y_1)^2\]

OpenStudy (anonymous):

but how the equation is for a parabola

OpenStudy (zarkon):

you have two points (2,-8) and (x,y^2) \[D^2=(x-2)^2+(y-(-8))^2\]

OpenStudy (zarkon):

oops (4x,y^2)

OpenStudy (anonymous):

it should be (x, square root of 4x)

OpenStudy (zarkon):

write \[x=\frac{y^2}{4}\] \[D^2=(x-2)^2+(y-(-8))^2=\left(\frac{y^2}{4}-2\right)^2+(y+8)^2\]

OpenStudy (zarkon):

differentiate with respect to y

OpenStudy (zarkon):

i get y=-4 and x=4

OpenStudy (anonymous):

 see the drawing with due recognition to quadrants we can see that focus lies (2,0) so perependicular distance being the shortest , we will have some co-ordinate on the line drwn which is of form (2,y) put this into y^2=4ax to get x and y it is (2,2root2)